A high resistance voltmeter is connected to a dry cell battery. The voltmeter reads 10 volts. When a 20 ohm resistor is connected across the terminals of the battery the voltmeter reading becomes 8 volts. What is the internal resistance of the battery?

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vidisha

If voltmeter resistance is high then it practically no current through it.

in first case it shows the emf, E of the cell

E=10 volts

in second case I=E/R+r  , R=20 and r is internal resistance

E-Ir=8

E-Er/(R+r) = 8

put value of E=10 and R=20

10 - 10r/20+r = 8

solving we get r=5 ohm

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively :Click here to download the toolbar..